a) Find the no-load value of vo in the circuit shown. b) Find vo when Rt is 150 kQ. c) How much power is dissipated in the 25㏀ Resisto r if the load terminals are accidentally short-circuited? d) What is the maximun power dissipated in the 75 ㏀ Resistor when R, is 150 ㏀? 25 kΩ 200 V RL


a) find the no-load value of v0 in the circuit shown
b) find v0 when RL is 150 k(ohms)
c) how much power is dissipated in the 25 k(ohms) resistor if the load terminals are accidentally short-circuited?
d) what is the maximum power dissipated in thr 75 k(ohms) resistor when RL is 150 k(ohms)?
Answer #1

A) no load value of Vo =200*75/(75+25) =150V

B)when RL =150kOhm

Rparallel = 75*150/(75+150)

= 50 kohm

Vo = 200*50/(50+25)

= 133.3 V

C) if short circuited, only resistance in the circuit will be 25kohm

Power = V^2/R = 200^2/25000 = 1.6 W

D) Maximum power = V^2/R = 133.3^2/75000

= 0.237 W